Looking at a RC circuit, using resistors and capacitors, calculate the RC time constant for discharging and charging capacitors. The voltage from the power supply is 10V. V = 0.63*V0, V = 0.37*V0
Answer #1

Dear student,

Find this solution, and RATE IT ,If you find it is helpful .your rating is very important to me.If any incorrectness ,kindly let me know I will rectify them soon.

Thanks for asking ..a) for charging RC - Circuit, V=V (1-7 %) = Here, t= time constant .=RC I at t=t, after one time constant, V= Voll-ef) = Vol1